Lauren is the youngest of four sisters. The average of their ages(all different whole numbers) is 9 years. What is the oldest that Lauren can be in years?
Answer by Alan3354(69443)

Lauren is the youngest of four sisters. The average of their ages(all different whole numbers) is 9 years. What is the oldest that Lauren can be in years?
------------
The max of Lauren's age --> the minimum of the others' ages, 1, 2 & 3
4*avg = 4*9 = 36
36 - (1+2+3) = 30
Lauren is 30 yrs of age, or younger.